If x+1/x = 10/3 then please find the value of x^3-1/x^3
 

Grade:9

1 Answers

Arun
5 years ago
Dear student
 
x + 1/x = 10/3
 
3x² -10x +3 = 0
 
x = 3 and 1/3
 
Now
 
x³ - 1/x³
If x = 3
 
= 27 - 1/27 = 728/27
 
Now if x = 1/3
 
1/27 - 27 = -728/27
 
 
Regards
Arun